# Physical partisan proximity outweighs online ties in predicting US voting outcomes

### v1

@@ Results
Finally, as a sensitivity analysis, we excluded the self-loops from the co-location and online connections networks. A self-loop represents the co-location or friendship probability between individuals in county $i$ and residents of the same county, leading to an edge between $i$ and itself. In the physical network, local exposure represents a substantial and fundamental aspect of an individual’s exposure, with a median value of counties’ extroversion equal to $0.27$. Conversely, while local exposure still plays a significant role in the online network, the network structure is more heterogeneous and characterized by higher external exposure (median value of extroversion equal to $1.19$). Detailed descriptions can be found in the SI Appendix (Fig. S3). Fig. S4 of the SI Appendix shows that when local exposure is excluded from the computation of $PE$, online ties outweigh physical proximity in predicting US political outcomes at the county level. Specifically, considering all the US counties, online exposure explains $65\%$ and $67\%$ ($R^{2}$) of the variance of voting patterns for Republicans and Democrats, respectively, compared to $61\%$ and $60\%$ explained by the physical proximity (Fig. S4a). However, predictive performances are lower when local exposure is not considered, particularly in the offline network, characterized by low heterogeneity and extroversion. Hence, local exposure appears to be a fundamental dimension in predicting voting outcomes.
Partisan segregation across demographic and socioeconomic factors
To investigate the levels of social segregation across political lines in the US, we estimate partisan segregation for each dimension as the net difference between exposure to Republicans and exposure to Democrats in a given county. The metric is scaled from 0 (indicating exclusive exposure to Democrats) to 1 (indicating exclusive exposure to Republicans). Intermediate values represent different mixing levels, with 0.5 indicating an equal share of exposure between the two electorates (see Materials and Methods Section for further details). We then investigate the relationship between partisan segregation and the county demographic and socioeconomic characteristics, focusing on metropolitan areas. To this aim, we predict partisan segregation for each network dimension using Gradient Boosting (GB) regressions, based on a set of demographic and socioeconomic factors: the shares of Hispanics and Latinos, African Americans, graduated individuals, unemployment, and urban population. Additionally, we enhance model explainability by computing SHAP values [48], which provide importance scores that represent the impact of each variable on the predicted outcome. We refer to the Materials and Methods for further details on machine learning models.
[…]
Figure 4: Partisan segregation across demographic and socioeconomic factors for each dimension. Maps and population-weighted distributions of the partisan segregation as captured by the Colocation Maps (a), Social Connectedness Index (b), and at the residential level (c). The red solid line represents the weighted average, while the black dashed line indicates balanced social mixing ($0.5$). For each dimension, the SHAP values distributions, computed from the GB regressions, highlight how the demographic and socioeconomic characteristics of the counties impact partisan segregation. Predictors are ordered by their impact on the final prediction, with points colored red for high and green for low values.
Overall, GB regressions effectively explain the variance of offline ($R^{2}=0.63$), online ($0.55$), and residential ($0.64$) segregation. Moreover, Fig. 4 shows the distribution of the SHAP values for each predictor related to offline, online, and residential partisan segregation in metropolitan areas. Predictors are ordered based on the impact on the prediction, with the most impactful variable at the top, and each point represents an observation. Finally, each point is colored based on its value, with red indicating high values and green indicating low values.
The share of graduated individuals is the best predictor for both partisan segregations computed from offline (average absolute impact equal to $0.06$) and online ($0.048$) networks. While counties with higher shares of graduated individuals tend to be exposed to Democrats (with higher Democratic segregation), counties with less educated individuals tend to be co-located and connected online with individuals who vote for the Republican party (Figs. 4a and 4b). In contrast, at the residential level, the educational level follows similar patterns but is the third predictor (mean average impact equal to $0.038$) in terms of variable importance (Fig. 4c).
Furthermore, partisan segregation is closely tied to the presence of ethnic communities. Specifically, the presence of a large share of African Americans is the second most relevant predictor for partisan segregation both online ($0.026$) and at the residential level ($0.042$) in metropolitan areas. Large communities of African Americans are associated with predominant exposure to Democrats, while areas with low African American presence tend to Republican segregation (Fig. 4). Finally, as shown in Fig. 4, the presence of Latinos and Hispanics is the least relevant predictor in all the dimensions, namely offline ($0.014$), online ($0.012$), and at the residential level ($0.016$). Interestingly, a larger share of Latinos and Hispanics in US counties are associated with higher levels of Democratic and Republican offline and online segregation.
@@ Discussion
In this study, we compared partisan exposure across physical and online spaces, investigated the relationship between partisan exposure and the voting patterns of US counties, and evaluated the association between partisan segregation in online and physical spaces and the underlying counties’ demographic and socioeconomic characteristics. Our findings highlight the dominant role of physical proximity and local exposure in predicting voting patterns in the United States, outperforming the influence of online ties and residential proximity. The result was consistent across both metropolitan and non-metro areas, as well as in swing and non-swing counties. Moreover, we found that partisan segregation is higher in offline social networks than online ones, and such difference is primarily driven by individuals’ educational attainment. The urban-rural divide significantly shapes offline and residential partisan segregation, with Republican voter segregation more prevalent in metro areas with large rural populations. Thus, physical partisan exposure, which is the best predictor of voting outcomes, is also strongly associated with educational attainment and urbanization levels.
Our findings contribute to a better understanding of the relationship between social networks, physical and online spaces, and political behavior. Previous studies have mainly focused on the impact of new media technologies, particularly online social media, on affective polarization and political elections [5, 6, 7, 8, 9, 10, 11, 12, 13], either through mechanisms of echo chambers and selective exposure [6, 49] or through interactions outside one’s local bubbles [11]. In this context, our study, by comparing physical and online networks, highlights the high relevance of physical space and offline ties in predicting political voting patterns at the county level, in both metro and non-metro areas, and especially where political voting is less predictable (i.e., swing counties). Local exposure and physical proximity, characterized by low heterogeneity of social exposure and low extroversion, not only account for social ties but also for visitation patterns to places where individuals spend their time, as captured by the co-location probabilities. Our findings call for a deeper investigation of geographic space, human behavior, and the social groups that share this space, which often tend to share beliefs and opinions through mechanisms of homophily [50].
@@
Moreover, partisan exposure and segregation vary across physical and online spaces, as well as between urban and rural areas. In line with the literature, our findings highlight that a large share of the US population, up to 10%, has low exposure to outgroup partisans, not only in the residential sphere [23], but also in physical social encounters, where experienced segregation is even greater than online partisan segregation. Both forms of experienced segregation are associated with individuals’ educational attainment, a known driver of socioeconomic segregation [51]. Furthermore, the tendency of Republican voter segregation in areas with large rural populations reveals differences between urban and rural areas related not only in terms of political party support [52] but also in partisan exposure [23]. In contrast, consistent with the literature, our findings show greater Democratic segregation in high-density urban areas [23]. These findings are relevant for public policy as they highlight how human activities, either online or in physical social encounters, reduce political social mixing, driven by educational and geographic disparities. Addressing these underlying factors can help policymakers to reduce social divisions and enhance social cohesion.
@@
Our study comes with some limitations. First, we compared online and physical networks using aggregated observational data from social media platforms. Despite the two-step reweighting applied to the Colocation Maps dataset [37], Facebook data might under-represent certain categories of people (e.g., young and older individuals) and the population of rural areas [53]. Moreover, these aggregated networks are considered at a rather coarse geographical level, the US county, whose population sizes vary widely, ranging from $83$ individuals (Loving County, TX) to over $10$ million inhabitants. Furthermore, we use the SCI as a proxy of online connections, but people interact through multiple social media platforms and have access to novel information from different sources. However, although some studies have attempted to map online interactions spatially [27, 54, 55], achieving comprehensive coverage across the entire US remains challenging. Finally, we acknowledge that co-location does not imply a face-to-face interaction between individuals. However, co-presence temporal networks have been found to have comparable structural and statistical features to face-to-face interaction temporal networks [56].
@@
Overall, our study underscores the centrality of physical space in understanding human and political behavior in both urban and rural areas. Despite challenges related to data limitations and privacy concerns, future studies on political polarization, partisan segregation, and their impact on political outcomes should not overlook the importance of real-world interactions in the physical space.
@@
We use the Social Connectedness Index (SCI) [33] dataset provided by Meta through its Data for Good program as a proxy of online connections. The SCI measures the social connectedness between administrative areas by computing the probability at which two random individuals, living respectively in administrative areas $i$ and $j$, are friends on Facebook, as of April 2016 [33]. The SCI is available at both the county and zip code levels. We analyze online connections at the county level to enable comparison with the Colocation Maps. Specifically, to be able to compare the online probabilities with the offline ones, we use the original version of the dataset adopted in Bailey et al. [33], which includes the relative probabilities of friendship between counties computed as follows:
$rel\_prob\_friend=C\frac{sci}{pop_{i}*pop_{j}},$ (1)
where $sci$ is the number of Facebook friendships between administrative regions $i$ and $j$ and $pop_{i}$ and $pop_{j}$ are the population size of $i$ and $j$. $C$ is a scaling factor, equal to $10^{12}$, that we drop to obtain the raw probabilities.
@@
We use the Colocation Maps [37] provided by Meta through its Data for Good program to measure offline exposure. Similarly to the SCI, the dataset provides the weekly co-location probability between two randomly selected individuals from two administrative areas $i$ and $j$. A co-location event (Fig. 1a) is registered when two individuals who enabled location-based services on the Facebook smartphone application are located in the same place (i.e., a level-16 Bing-tile, approximately 600 by 600 meters at the equator) at the same time for at least 5 minutes. The co-location probability between geographic regions $i$ and $j$ is then determined by the number of co-location events between individuals from $i$ and $j$ weighted by their population sizes. As described by Iyer et al. [37], the authors applied a two-steps reweighting to enhance representativeness. Specifically, the first step aligns the Facebook user data to the general population, and the second step reweights the Colocation Maps population to the Facebook user base. The dataset has been designed and used mainly for epidemiological purposes [57, 58]. For the aim of our study, we collected weekly co-location probabilities between US counties from August 2022 to January 2024, for a total of $76$ weeks/observations. We then use the average co-location value over the full period to mitigate seasonality effects and data sparsity and reduce computational errors in data measurement, ensuring a more realistic measure of physical exposure between counties.
@@
We compare online and offline networks with a measure of residential proximity, based on individuals’ residential locations. To this purpose, we use the dataset provided by R. Enos [38] supporting the recent study by Brown et al. [23], which includes the conditional probability of exposure to Democratic and Republican voters, conditioned on an individual’s political affiliation (either Democrat or Republican). The probabilities are computed based on the nearest 1,000 individuals, but the results remain consistent even considering a larger sample size of 50,000 individuals. The dataset is derived from voter registration records, which require individuals in the majority of US states to declare their party affiliation and it provides information at different geographical levels. To have a measure of partisan exposure at the residential level, we compute the probabilities of exposure to Democrats and Republicans for a random individual for each US county. Specifically, from the available raw data, we compute the partisan exposure to Republicans for a county $i$ as follows:
@@
We define the voting patterns of the counties in the contiguous United States following the normal vote concept [39] to account for candidate- and election-specific influences on voting outcomes. Specifically, the share of Republican and Democratic voters for each county is computed as the average of the last three presidential elections (i.e., 2012, 2016, and 2020). We obtained data on presidential election results from the work conducted by Algara and Amlani [59].
@@
To evaluate the relationship between partisan segregation and county characteristics, we obtain data from the 2017-2021 5-year estimates of the American Community Survey (ACS) provided by the United States Census Bureau [60]. Specifically, we use data related to the shares of Hispanics and Latinos, African Americans, unemployed individuals, and urban population. Moreover, we use the information about the share of graduated individuals for each county provided by Chetty et al. [61] through the Opportunity Insights repository [62]. We select this information from a broader set of counties’ characteristics by computing Variance Inflation Factors (VIFs) to address multicollinearity issues (See SI for further details) and drawing on the literature on the determinants of county outcomes in US presidential elections. Specifically, ethnicity is one of the strongest predictors of political outcomes, with African Americans tending to vote for the Democratic party [42, 46]. Additionally, educational attainment, which is highly correlated with income (addressed through VIF), is associated with Democratic voting tendencies [42, 44]. Finally, unemployment is one of the economic indicators that impact political outcomes [45], as well as the urban-rural characteristics and divide [43, 52].

### v2

@@ Results
Finally, as a sensitivity analysis, we excluded the self-loops from the co-location and online connections networks. A self-loop represents the co-location or friendship probability between individuals in county $i$ and residents of the same county, leading to an edge between $i$ and itself. In the physical network, local exposure represents a substantial and fundamental aspect of an individual’s exposure, with a median value of counties’ extroversion equal to $0.27$. In contrast, while local exposure still plays a significant role in the online network, the network structure is more heterogeneous and characterized by higher external exposure (median value of extroversion equal to $1.19$). Detailed descriptions can be found in the Supplementary Information (Fig. S1). Fig. S5 of the Supplementary Information shows that when local exposure is excluded from the calculation of $PE$, online ties outweigh physical proximity in predicting US political outcomes at the county level. Specifically, considering all the US counties, online exposure explains $65\%$ and $67\%$ ($R^{2}$) of the variance of voting patterns for Republicans and Democrats, respectively, compared to $61\%$ and $60\%$ explained by the physical proximity (Fig. S5a). However, predictive performance is lower when local exposure is not considered, particularly in the offline network, characterized by low heterogeneity and extroversion. Hence, local exposure appears to be a fundamental dimension in predicting voting outcomes.
Offline partisan exposure better predicts individuals’ voting behavior
Figure 4: Average marginal effects of partisan exposure (online and offline) on vote choice. Offline partisan exposure has a stronger average marginal effect on vote choice than online partisan exposure, both in terms of exposure to Democrats (a) and Republicans (b). Logit models control for respondents’ age, ethnicity, educational attainment, and place of residence (metro or non-metro area). The dependent variable is binary, with 0 for voting Democrat and 1 for voting Republican.
We combine large-scale analysis with individual survey data to better understand the relationship between partisan exposure in physical and digital spaces and vote choice. To this end, we leverage the 2020-2022 Social Media Study dataset [60] provided by the American National Election Studies (ANES), which consists of an online survey panel conducted in three waves (pre- and post-election 2020, and the 2022 midterm elections). As explained in the Materials and Methods Section, our focus is on the first two waves, including all respondents who declared their vote choice in the 2020 presidential election for either the Democratic or Republican party, had a valid vote, and responded to the following questions: "How many of your friends and family are Democrats?", "How many of your friends and family are Republicans?", "How many of your Facebook friends are Democrats?", and "How many of your Facebook friends are Republicans?". The final sample consists of 2420 respondents. Descriptive information about the variables can be found in the Supplementary Information (Table S26).
As described in the Materials and Methods section, we performed four logit models to distinguish between exposure to Democrats and Republicans and to account for each wave (pre- and post-election). The dependent variable is vote choice, encoded as a binary variable with 0 for the Democratic party and 1 for the Republican party. We control for the age, ethnicity, educational attainment and place of residence of the respondents (metro or non-metro area), and compute and apply post-stratification weights based on gender, age, education and ethnicity using the 5-year estimates 2020-2024 of the American Community Survey (ACS). Detailed results can be found in the Supplementary Information (S27).
Fig. 4 shows the Average Marginal Effects (AME) of offline and online partisan exposure on vote choice across the two waves. As shown, offline exposure to Democrats has a stronger effect on voting for the Democratic party than online exposure (4a), while offline exposure to Republicans more significantly impacts voting for the Republican party (4b). Specifically, in the first wave, the average marginal effect of the exposure to Democrats is 0.163-0.163 offline compared to 0.061-0.061 online, while in the second wave it is 0.159-0.159 compared to 0.054-0.054. Similarly, related to the exposure to Republicans, the average marginal effect is $0.127$ offline compared to $0.089$ online in the first wave, and $0.174$ compared to $0.045$ in the second wave.
@@ Discussion
In this study, we compared partisan exposure across physical and online spaces, evaluated the association between partisan segregation in online and physical spaces and the underlying counties’ demographic and socioeconomic characteristics, and investigated the relationship between partisan exposure and vote choice at both the individual and county levels. Our findings highlight the dominant role of experienced physical proximity and local exposure in predicting vote choice in the United States at both the individual and county levels, outperforming the influence of online ties and residential proximity. The result was consistent across both metropolitan and non-metro areas. Moreover, we found that partisan segregation is higher in offline social networks than online ones, and such a difference is primarily driven by individuals’ educational attainment. The urban-rural divide significantly shapes offline and residential partisan segregation, with Republican voter segregation more prevalent in metro areas with large rural populations. Thus, physical partisan exposure, which is the best predictor of voting outcomes, is also strongly associated with educational attainment and urbanization levels. From this perspective, our findings contribute to a better understanding of the relationship between social networks, physical and online spaces, and political behavior, highlighting the effects of partisan segregation on voting polarization, with potential implications for affective polarization [4].
Despite significant attention to the impact of new media technologies, particularly online social media, on affective polarization and political elections [6, 7, 8, 9, 10, 11, 12, 13, 14], either through mechanisms of echo chambers and selective exposure [7, 25] or through interactions outside of one’s local bubbles [12], our findings highlight the greater relevance of physical space and offline ties in predicting political behavior. This is demonstrated through both large-scale (i.e., county-level) and individual survey analyses, enhancing our understanding of the relationship between partisan exposure and political vote. From this perspective, offline social networks, local exposure and physical proximity, characterized by low heterogeneity of social exposure and low extroversion, better reflect vote choice at the individual and county levels compared to online ties. Similar to prior research [53, 54], our results suggest that even brief, passive, and casual forms of exposure, which can be captured by co-location events, can influence or be associated with political behavior. Therefore, beyond the form of exposure, context matters, with offline contexts outweighing online exposure. In line with the literature that examines the association between social networks and vote choice [71, 50], these findings confirm the link between homogeneous social networks and political vote, particularly in physical space, where Americans experience greater partisan segregation. This underscores the persistent influence of offline social ties compared to online ones, despite increasing attention to social media platforms and their potential impact on vote choice and affective polarization.
@@
In general, partisan exposure and segregation vary between physical and online spaces, as well as between urban and rural areas. In line with the literature, our findings highlight that up to 10% of the American population experiences a low exposure to outgroup partisans, not only in the residential sphere [17], but also in physical social encounters. Our analysis shows that experienced segregation is even greater than online partisan segregation, consistent with previous studies [26, 13, 27]. Both forms of experienced segregation are associated with the educational attainment of individuals, a known driver of socioeconomic segregation [72]. Furthermore, the tendency of Republican voter segregation in areas with large rural populations reveals differences between urban and rural areas related not only in terms of political party support [73], but also in partisan exposure [17]. In contrast, consistent with the literature, our findings show greater Democratic segregation in high-density urban areas [17].
@@
Our study comes with some limitations. First, we compared online and physical networks using aggregated observational data from social media platforms. Despite the two-step reweighting applied to Colocation Maps [57], the Facebook data might underrepresent certain categories of people (e.g., young and old individuals) and the population of rural areas [74]. Moreover, these aggregated networks are considered at a rather coarse geographical level, the US county, whose population sizes vary widely, ranging from $83$ individuals (Loving County, TX) to over $10$ million inhabitants. Furthermore, we use the SCI as a proxy of online connections, but people interact through multiple social media platforms and have access to novel information from different sources. However, although some studies have attempted to spatially map online interactions [33, 75, 76], achieving comprehensive coverage throughout the US remains challenging. Moreover, we acknowledge that co-location does not imply a face-to-face interaction between individuals. However, co-presence temporal networks have been found to have structural and statistical features comparable to face-to-face interaction temporal networks [77]. Additionally, individuals’ partisan exposure, as measured by the ANES Social Media Study both before and after the 2020 US presidential election, reflects self-perceived partisan exposure and may be subject to self-report biases, such as social desirability or recall inaccuracy. Finally, as mentioned, these datasets capture different types of contact, varying in duration, selectivity, and the extent to which they are active or passive. This is also reflected in the mismatch between county- and individual-level analyses, with the latter capturing more active forms of exposure, typically driven by friendship or kinship ties. Future work should aim at measuring the partisan balance of individual’s social networks, online and offline, based on one comprehensive data source.
@@
In conclusion, our study underscores the centrality of physical space in understanding human and political behavior, with potential implications for affective and voting polarization driven by partisan segregation. Despite challenges related to data limitations and privacy concerns, future studies on political polarization, social networks, and partisan segregation, and their impact on political outcomes, should not overlook the importance of real-world interactions in physical space.
@@
We use the Social Connectedness Index (SCI) [58] dataset provided by Meta through its Data for Good program as a proxy of online connections. The SCI measures the social connectedness between administrative areas by computing the probability at which two random individuals, living respectively in administrative areas $i$ and $j$, are friends on Facebook, as of April 2016 [58]. The SCI is available at the county and zip code levels. We analyze online connections at the county level to enable comparison with the Colocation Maps. Specifically, to be able to compare the online probabilities with the offline ones, we use the original version of the dataset adopted in Bailey et al. [58], which includes the relative probabilities of friendship between counties computed as follows:
$rel\_prob\_friend=C\frac{f_{ij}}{pop_{i}*pop_{j}},$ (1)
where $f_{ij}$ is the number of Facebook friendships between administrative regions $i$ and $j$ and $pop_{i}$ and $pop_{j}$ are the population size of $i$ and $j$. $C$ is a scaling factor, equal to $10^{12}$, which we drop to obtain the raw probabilities.
@@
We use the Colocation Maps [57] provided by Meta through its Data for Good program to measure offline exposure. Similarly to the SCI, the dataset provides the weekly co-location probability between two randomly selected individuals from two administrative areas $i$ and $j$. A co-location event (Fig. 1a) is registered when two individuals who enabled location-based services on the Facebook smartphone application are located in the same place (i.e., a level-16 Bing-tile, approximately 600 by 600 meters at the Equator) at the same time for at least 5 minutes. The co-location probability between geographic regions $i$ and $j$ is then determined by the number of co-location events between individuals from $i$ and $j$ weighted by their population sizes. As described by Iyer et al. [57], the authors applied a two-step reweighting to enhance representativeness. Specifically, the first step aligns the Facebook user data with the general population, and the second step reweights the Colocation Maps population to the Facebook user base. The dataset has been designed and used mainly for epidemiological purposes [78, 79]. For the aim of our study, we collected weekly co-location probabilities between US counties from August 2022 to January 2024, for a total of $76$ weeks/observations. We then use the average co-location value over the full period to mitigate seasonality effects and data sparsity and reduce computational errors in data measurement, ensuring a more realistic measure of physical exposure between counties.
@@
We compare online and offline networks with a measure of residential proximity, based on the residential locations of individuals. To this purpose, we use the dataset provided by R. Enos [59] supporting the recent study by Brown et al. [17], which includes the conditional probability of exposure to Democratic and Republican voters, conditioned on an individual’s political affiliation (either Democrat or Republican). The probabilities are computed based on the nearest 1,000 individuals, but the results remain consistent even considering a larger sample size of 50,000 individuals. The dataset is derived from voter registration records, which require individuals in the majority of US states to declare their party affiliation and it provides information at different geographical levels. To have a measure of partisan exposure at the residential level, we compute the probabilities of exposure to Democrats and Republicans for a random individual for each US county. Specifically, from the available raw data, we compute the partisan exposure to Republicans for a county $i$ as follows:
@@
ANES 2020-2022 Social Media Study
We analyze individual-level exposure to Democrats and Republicans and their relationship with political voting in the 2020 presidential election using the 2020-2022 Social Media Study provided by the American National Election Studies (ANES) [60]. The study is based on an online survey panel of 3 waves (that is, before and after the 2020 elections, and the 2022 midterm elections). We focus on the first two waves, which include 5,750 and 5,277 respondents, respectively. From this sample, we consider only those respondents who declared their political preference in the second wave ($w2presvtwho$), have valid voting records (vote20_matchvote20\_match), and answered all the questions of interest. The final sample consists of 2420 respondents. Specifically, we focus on the following questions: “How many of your friends and family are Democrats?", “How many of your friends and family are Republicans?",“How many of your Facebook friends are Democrats?" and “How many of your Facebook friends are Republicans?" Each of these questions is measured on a scale from 1 to 5, where 1 indicates “None or almost none", and 5 indicates "All or nearly all". We then compute new post-stratification weights based on gender, age, education, and ethnicity using the 2020-2024 5-year estimates of the American Community Survey (ACS).
@@
We define the voting patterns of counties in the contiguous United States following the normal vote concept [61] to account for candidate- and election-specific influences on voting outcomes. Specifically, the share of Republican and Democratic voters for each county is computed as the average of three presidential elections (i.e., 2012, 2016, and 2020). We obtained data on presidential election results from the work by Algara and Amlani [80].
@@
To evaluate the relationship between partisan segregation and county characteristics, we obtain data from 5-year estimates of the American Community Survey (ACS) for 2017-2021 provided by the US Census Bureau [81]. Specifically, we used data related to the share of Hispanics and Latinos, African Americans, unemployed individuals, and urban population. In addition, we use information about the share of graduated individuals for each county provided by Chetty et al. [82] through the Opportunity Insights repository [83]. We select this information from a broader set of counties’ characteristics by computing Variance Inflation Factors (VIFs) to address multicollinearity issues and drawing on the literature on the determinants of county outcomes in US presidential elections. Specifically, ethnicity is one of the strongest predictors of political outcomes, and African Americans tend to vote for the Democratic party [65, 69]. Furthermore, educational attainment, which is highly correlated with income (addressed through VIF), is associated with Democratic voting tendencies [65, 67]. Finally, unemployment is one of the economic indicators that affect political outcomes [68], as well as the characteristics and divide of urban-rural regions [66, 73].